Q=180√LK^0.8

L=5

K=10

1).Construct marginal product of labour and marginal product of capital

2If L=5, K=10 Measure average product of labour and average product of capital

3.return to scale production function

4.if K=5, L=10 calculate marginal rate of products.

5.law of diminishing marginal returns of labour and law of diminishing marginal returns of capital

6.how about the cross partial effect of this production function


1
Expert's answer
2021-08-04T09:38:27-0400

"Q=180\sqrt{L}K^{0.8}\\L=5\space K=10"


1.

"MPL=\frac{dQ}{dL}=\frac{180}{2\sqrt L}K^{0.8}=\frac{90}{\sqrt L}K^{0.8}"


"MPK=\frac{dQ}{dK}=180\times0.8\sqrt{L}\space K^{-0.2}\\MPK=144\sqrt{L}K^{-0.2}"


2.

"Q=180\sqrt{L}K^{0.8}"


"\\APL=\frac{Q}{L}=\frac{180\sqrt{L}K^{0.8}}{L}=\frac{180}{\sqrt{L}}K^{0.8}"


"\\APL=\frac{180(10)^{0.8}}{5}=80.49845(10)^{0.8}"


"\\APL=507.9"


"\\APK=\frac{Q}{K}=\frac{180\sqrt{L}K^{0.8}}{K}"


"\\APK=\frac{180\sqrt{L}}{K^{0.2}}=\frac{180\sqrt{5}}{10^{0.2}}=\frac{402.49}{10^{0.2}}"


"\\\\APK=\frac{402.49}{1.5848932}=253.9554"


3.

"Q(K,L)=180\sqrt{L}K^{0.8}"

Scaling both factors of production by t>t

"Q(tK,tL)=180\sqrt{tL}(tK)^{0.8}\\"


"=180(t)^{\frac{1}{2}}\sqrt{L}(t)^{0.8}(K)^{0.8}\\=180\sqrt{L}(K)^{0.8}(t)^{0.8}\\=180\sqrt{L}(K)^{0.8}(t)^{0.5+0.8}\\=(t)^{1.3}180\sqrt{L}(K)^{0.8}"

"Q(tK,tL)=Q(K,L)(t)^{1.3}"


"(t)^{1.3}>(t)" implies increasing returns to scale

4.

"MRTS=\frac{MPL}{MPK}=\frac{90}{\sqrt{L}}(K)^{0.8}\times\frac{1}{144\sqrt{L}(K)^{-0.2}}"


"=\frac{90}{144} \frac{K^{0.8}(K)^{0.2}}{L}"


"=\frac{90}{144}\frac{K}{L}"


"=\frac{10}{16}\frac{K}{L}"


"MRTS=\frac{5}{8}\frac{K}{L}"


5.

Cross partial effects

"MPL=\frac{90}{\sqrt{L}}(K)^{0.8}"


"\frac{dMPL}{dK}=90\frac{(0.8)}{\sqrt{L}}(K)^{-0.2}\\=\frac{72(K)^{-0.2}}{\sqrt{L}}>0"


"MPK=144\sqrt{L}(K)^{-0.2}"


"\frac{dMPK}{dL}=\frac{144}{2\sqrt{L}}(K)^{-0.2}=\frac{72}{\sqrt{L}}(K)^{-0.2}>0"